OpenStudy (anonymous):

Simplify: (y-(4/y))/(3+(6/y))

OpenStudy (anonymous):

\[\frac{y-\frac{4}{y}}{3+\frac{6}{y}}\] like this? multiply top and bottom by y to clear the fraction in one step

OpenStudy (anonymous):

\[\frac{y-\frac{4}{y}}{3+\frac{6}{y}}\times \frac{y}{y}=\frac{y^2-4}{3y+6}\]

OpenStudy (anonymous):

yes, like that. the answer i know is (y-2)/3

OpenStudy (anonymous):

then factor the numerator and denominator to get \[\frac{(y+2)(y-2)}{3(y+2)}=\frac{y-2}{3}\]

OpenStudy (anonymous):

Thank you so much!!! I appreciate it.